OpenStudy (anonymous):

Solve for d. d - 2 d - 3 = 1 -m A) d = m + 3 m + 1 B) d = 2m + 3 m + 1 C) d = 2m + 1 m + 3 D) d = 2m + 3 m - 1

DivineSolar (divinesolar):

Which one are we solving? both?

OpenStudy (anonymous):

\[\frac{ d-2}{ d-3 }=\frac{ 1 }{ -m }\]

OpenStudy (anonymous):

hello are you there

DivineSolar (divinesolar):

Sorry

DivineSolar (divinesolar):

Yes i am here.

OpenStudy (anonymous):

can u solve it

DivineSolar (divinesolar):

Here you go. Hope my explanation through the work is good enough for you to find your answer.

OpenStudy (anonymous):

thanks

DivineSolar (divinesolar):

Sure thing.